# Quantum Channel Testing in Average-Case Distance

### v1

@@ 6.1 Upper bounds for arbitrary channels
In this subsection we give an ancilla-free, non-adaptive upper bound for testing identity to an arbitrary channel, measuring the distance between two channels by the $\ell_{2}$ distance between their Choi states. By relating this distance to ACID and diamond distances, we then obtain upper bounds for channel certification in ACID and diamond distances too.
We begin with the following lemma:
Lemma 6.1.
For all matrices $X,Y\in\mathbb{C}^{d\times d}$,
@@
Gu [Gu13] proved a significant generalization of Lemma 6.1 using Weingarten calculus. For completeness and simplicity, below we present a self-contained proof of Lemma 6.1 using only Eq. 5. The following argument was essentially made by Montanaro and de Wolf [MdW13, proof of Proposition 21].
Proof.
We have
where the last equality is by Eq. 5. Clearly
and furthermore
The result follows by combining the above three equations. ∎
Using Lemma 6.1 we prove the following equality:
Lemma 6.2.
where $\boldsymbol{\psi}\in\mathsf{D}(d)$ is Haar random.
We remark that Fawzi et al. proved Lemma 6.2 in the case where $\mathcal{L}$ is the identity and $\mathcal{K}$ is an arbitrary channel [FFGO23, Lemma A.1], as well as in the case where $\mathcal{L}=\mathcal{K}$ is the difference between an arbitrary channel and the completely depolarizing channel [FFGO23, Lemma C.2].
Proof.
We will use the following lemma to bound the success probability of our channel certification algorithm:
Lemma 6.3.
Let $\mathcal{L}$ be the difference between two channels, and let $0<\varepsilon\leq{\|J_{\mathcal{L}}\|}_{2}$. Then there exists a positive integer $k\leq\mathopen{}\mathclose{{}\left\lceil\log\mathopen{}\mathclose{{}\left(4/\varepsilon^{2}}\right)}\right\rceil$ such that for Haar random $\boldsymbol{\psi}$,
$\mathrm{Pr}\mathopen{}\mathclose{{}\left(\frac{\varepsilon^{2}}{4}2^{k-1}<{\|\mathcal{L}(\boldsymbol{\psi})\|}_{2}^{2}\leq\frac{\varepsilon^{2}}{4}2^{k}}\right)\geq\frac{2^{-k}}{\mathopen{}\mathclose{{}\left\lceil\log\mathopen{}\mathclose{{}\left(4/\varepsilon^{2}}\right)}\right\rceil}.$
More generally, Lemma 6.3 holds for any Hermitian-preserving superoperator $\mathcal{L}$ satisfying ${\|\mathcal{L}\|}_{1}\leq 2$, a criterion which holds for the difference between two channels by the triangle inequality.
Proof.
Let $\boldsymbol{X}={\|\mathcal{L}(\boldsymbol{\psi})\|}_{2}^{2}$. Write $\mathcal{L}\in\mathsf{S}(d,*)$, and observe that by the case of Lemma 6.2 where $\mathcal{K}=\mathcal{L}$,
$\varepsilon^{2}\leq{\|J_{\mathcal{L}}\|}_{2}^{2}\leq{\|J_{\mathcal{L}}\|}_{2}^{2}+{\|\mathcal{L}(I/d)\|}_{2}^{2}=\frac{d+1}{d}\operatorname*{\mathbb{E}}[\boldsymbol{X}]\leq 2\operatorname*{\mathbb{E}}[\boldsymbol{X}].$ (23)
Let $t=\mathopen{}\mathclose{{}\left\lceil\log\mathopen{}\mathclose{{}\left(4/\varepsilon^{2}}\right)}\right\rceil$, and define disjoint intervals
Since
$0\leq\boldsymbol{X}\leq{\|\mathcal{L}(\boldsymbol{\psi})\|}_{1}^{2}\leq{\|\mathcal{L}\|}_{1}^{2}\leq 4\leq\varepsilon^{2}2^{t}$
pointwise, there exists a unique $k$ such that $\boldsymbol{X}$ is in $P_{k}$, and therefore
Now we give our upper bound for channel certification in $\ell_{2}$ distance between Choi states:
Theorem 6.4.
@@
Proof.
[…]
$\sum_{k=1}^{t}2^{k+1}t\cdot\delta=\mathopen{}\mathclose{{}\left(2^{t+2}-4}\right)\varepsilon^{2}/96\leq 2^{\log(4/\varepsilon^{2})+3}\varepsilon^{2}/96=1/3.$
If ${\|J_{\mathcal{M}}-J_{\mathcal{N}}\|}_{2}\geq\varepsilon$, then by Lemma 6.3 there exists some fixed $k\in[t]$ such that
$\mathrm{Pr}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\|\mathcal{M}(\boldsymbol{\psi})-\mathcal{N}(\boldsymbol{\psi})}\right\|_{2}^{2}>\varepsilon^{2}2^{k-3}}\right)\geq 2^{-k}/t,$
@@
$\mathopen{}\mathclose{{}\left(1-(1-\delta)2^{-k}/t}\right)^{2^{k+1}t}\leq\exp\mathopen{}\mathclose{{}\left(-(1-\delta)2^{-k}/t\cdot 2^{k+1}t}\right)=\exp\mathopen{}\mathclose{{}\left(-(1-\delta)2}\right)\leq\exp(-4/3)<1/3.\qed$
@@
Let $v(P)$ denote the total variation distance between the uniform distribution and a probability distribution $P$ with finite support. Paninski [Pan08] gave an algorithm $\textsc{TestMixed}\mathopen{}\mathclose{{}\left(\delta,d}\right)$ that takes as input $O(\sqrt{d}/\delta^{2})$ samples from a probability distribution $P$ on $[d]$, and decides whether $P$ is the uniform distribution or $v(P)\geq\delta$ with success probability at least $2/3$. We may assume without loss of generality that the success probability of $\textsc{TestMixed}\mathopen{}\mathclose{{}\left(\delta,d}\right)$ is at least $1-1/30000$, by repetition and majority vote. For a universal constant $c$ and all $d\in\mathbb{N}$, Fawzi et al. [FFGO23, Lemma E.1] gave a random $cd$-outcome POVM $\boldsymbol{P}_{d}$ on $\mathbb{C}^{d\times d}$ such that for all fixed density matrices $\rho\in\mathsf{D}(d)$, if $\boldsymbol{P}_{d}(\rho)$ denotes the probability distribution defined by performing $\boldsymbol{P}_{d}$ on $\rho$, then
@@
Now consider the case where ${\|\mathcal{M}-\mathcal{N}\|}_{J}\geq\varepsilon$. Let $\boldsymbol{\psi}\in\mathsf{D}(d_{\mathrm{in}})$ be the Haar random state from Algorithm 2, and let $\boldsymbol{X}=\mathopen{}\mathclose{{}\left\|\mathcal{M}(\boldsymbol{\psi})-I/d_{\mathrm{out}}}\right\|_{2}^{2}$. Fawzi et al. [FFGO23, middle of Page 40] proved that $\mathrm{Pr}\mathopen{}\mathclose{{}\left(\boldsymbol{X}\geq\operatorname*{\mathbb{E}}[\boldsymbol{X}]/2}\right)\geq 1/1000$. If $\boldsymbol{X}\geq\operatorname*{\mathbb{E}}[\boldsymbol{X}]/2$ and $v(\boldsymbol{P}_{d_{\mathrm{out}}}(\mathcal{M}(\boldsymbol{\psi})))\geq\sqrt{\boldsymbol{X}}/20$—an event which occurs with probability at least $1/2000$, by the definition of $\boldsymbol{P}_{d_{\mathrm{out}}}$—then
$\varepsilon$ $\leq{\|J_{\mathcal{M}}-I/d_{\mathrm{in}}d_{\mathrm{out}}\|}_{1}$
$\leq\sqrt{2(d_{\mathrm{in}}+1)d_{\mathrm{out}}\boldsymbol{X}}$
@@ 6.2 Upper bounds for erasure, unitary, and pure state replacement channels
Proof sketch.
This is a special case of Lemma 6.2. ∎
Now we reduce the task of testing identity to $\mathcal{I}\otimes\mathcal{N}$ to that of testing identity to $\mathcal{N}$, for an arbitrary channel $\mathcal{N}$. To match the context in which we will apply this reduction, we phrase it in terms of ancilla-free, non-adaptive channel testers with perfect completeness, but the proof can easily be adopted to other query models (recall Section 2.3) and to channel testers with imperfect completeness as well.
@@
and queries to $\mathcal{P}$ can be simulated using queries to $\mathcal{M}$. The identity channel is just an erasure channel with input dimension equal to the output dimension, so the result follows by the above argument.
@@ 6.3 Lower bound for the completely depolarizing channel
We use the well-known fact that
$d_{\mathrm{TV}}(P,Q)\leq\sqrt{2}\cdot d_{\mathrm{H}}(P,Q)$
for all distributions $P$ and $Q$, where
$d_{\mathrm{H}}(P,Q)=\sqrt{1-\sum_{x}\sqrt{P(x)Q(x)}}$
denotes Hellinger distance. Let
$P=\bigotimes_{i=1}^{n}P_{i},$ $Q=U^{\otimes n},$
and write
@@
which rearranges to [^24] In fact, the stronger inequality $\sqrt{1-x}+\sqrt{1+x}\geq 2-(2-\sqrt{2})x^{2}$ holds, but proving this is more time-consuming and is not necessary for our purposes.
$\sqrt{1-x}+\sqrt{1+x}\geq 2-x^{2}.$
It follows that
@@
Recalling 2.3 of ancilla-free, non-adaptive channel testers, write
@@ 6.5 Tomography
$\mathopen{}\mathclose{{}\left\|\sigma_{n}-\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[|\boldsymbol{\mathcal{N}}\rangle\!\langle\boldsymbol{\mathcal{N}}|^{\otimes 2}}\right]}\right\|_{1}=\mathopen{}\mathclose{{}\left\|\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[|\boldsymbol{\mathcal{N}}\rangle\!\langle\boldsymbol{\mathcal{N}}|}\right]\otimes\mathopen{}\mathclose{{}\left(\rho_{\boldsymbol{\mathcal{N},q}}-|\boldsymbol{\mathcal{N}}\rangle\!\langle\boldsymbol{\mathcal{N}}|}\right)}\right\|_{1}=\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left\|\rho_{\boldsymbol{\mathcal{N},n}}-|\boldsymbol{\mathcal{N}}\rangle\!\langle\boldsymbol{\mathcal{N}}|}\right\|_{1}\leq 0.2,$
@@
$\leq$ $\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[(S_{\boldsymbol{\mathcal{N}}}(\mathsf{B})+S_{\boldsymbol{\mathcal{N}}}(\mathsf{C}))-(S(\mathsf{C})-S(\mathsf{B}))-(S_{\boldsymbol{\mathcal{N}}}(\mathsf{C})-S_{\boldsymbol{\mathcal{N}}}(\mathsf{B}^{\prime}))+(S(\mathsf{B}^{\prime})+S(\mathsf{C}))}\right]$ Eqs. 12 and 13
@@ Acknowledgments
GR, AD, TG are supported by the EPSRC New Horizons grant EP/X018180/1. SS is supported by a Royal Commission for the Exhibition of 1851 Research Fellowship. TG and HA are supported by ERC Starting Grant 101163189 and UKRI Future Leaders Fellowship MR/X023583/1. We thank Min-Hsiu Hsieh, Tony Metger, Jon Wright, and Henry Yuen for helpful discussions.

### v2

@@ 6.1 Upper bounds for arbitrary channels
We first prove the following:
Theorem 6.1 (Channel certification in $\ell_{2}$ distance between Choi states).
The proof uses the following result of Bao and Yao [Bao23Junta], of which we provide a (somewhat different) proof in Appendix B for completeness:
Lemma 6.2 ([Bao23Junta, Proposition 15]).
If $\mathcal{L}\in\mathsf{S}(d,*)$ is the difference between two channels, then
$\frac{d+1}{d}\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[{\|\mathcal{L}(\boldsymbol{\psi})\|}_{2}^{2}}\right]={\|J_{\mathcal{L}}\|}_{2}^{2}+{\|\mathcal{L}(I/d)\|}_{2}^{2}$
@@
Proof of Theorem 6.1.
[…]
$\sum_{k=1}^{t}2^{k+1}t\cdot\delta=\mathopen{}\mathclose{{}\left(2^{t+2}-4}\right)\varepsilon^{2}/384\leq 2^{\log(1/\varepsilon^{2})+7}\varepsilon^{2}/384=1/3.$
@@
Now suppose ${\|J_{\mathcal{M}}-J_{\mathcal{N}}\|}_{2}\geq\varepsilon$. Below we prove that there exists some fixed $k\in[t]$ such that
$\mathrm{Pr}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\|\mathcal{M}(\boldsymbol{\psi})-\mathcal{N}(\boldsymbol{\psi})}\right\|_{2}^{2}>\varepsilon^{2}2^{k-3}}\right)\geq 2^{-k}/t,$ (22)
@@
$\mathopen{}\mathclose{{}\left(1-(1-\delta)2^{-k}/t}\right)^{2^{k+1}t}\leq\exp\mathopen{}\mathclose{{}\left(-(1-\delta)2^{-k}/t\cdot 2^{k+1}t}\right)=\exp\mathopen{}\mathclose{{}\left(-(1-\delta)2}\right)\leq\exp(-4/3)<1/3.$
Toward establishing Eq. 22 for some $k\in[t]$, let $\mathcal{L}=\mathcal{M}-\mathcal{N}$ and $\boldsymbol{X}={\|\mathcal{L}(\boldsymbol{\psi})\|}_{2}^{2}$. By Lemma 6.2,
$\varepsilon^{2}\leq{\|J_{\mathcal{L}}\|}_{2}^{2}\leq{\|J_{\mathcal{L}}\|}_{2}^{2}+{\|\mathcal{L}(I/d_{\mathrm{in}})\|}_{2}^{2}=\frac{d_{\mathrm{in}}+1}{d_{\mathrm{in}}}\operatorname*{\mathbb{E}}[\boldsymbol{X}]\leq 2\operatorname*{\mathbb{E}}[\boldsymbol{X}].$ (23)
Define disjoint intervals
By the triangle inequality ${\|\mathcal{L}\|}_{1}\leq{\|\mathcal{M}\|}_{1}+{\|\mathcal{N}\|}_{1}=2$, so
$0\leq\boldsymbol{X}\leq{\|\mathcal{L}(\boldsymbol{\psi})\|}_{1}^{2}\leq{\|\mathcal{L}\|}_{1}^{2}\leq 4\leq\varepsilon^{2}/4\cdot 2^{t}$
pointwise, so there exists a unique $k$ such that $\boldsymbol{X}$ is in $P_{k}$, and therefore
@@
For a probability distribution $P$ with finite support, let $v(P)$ denote the total variation distance between $P$ and the uniform distribution. Paninski [Pan08] gave an algorithm $\textsc{TestMixed}\mathopen{}\mathclose{{}\left(\delta,d}\right)$ that takes as input $O(\sqrt{d}/\delta^{2})$ samples from a probability distribution $P$ on $[d]$, and decides whether $P$ is the uniform distribution or $v(P)\geq\delta$ with success probability at least $2/3$. We may assume without loss of generality that the success probability of $\textsc{TestMixed}\mathopen{}\mathclose{{}\left(\delta,d}\right)$ is at least $1-1/30000$, by repetition and majority vote. For a universal constant $c$ and all $d\in\mathbb{N}$, Fawzi et al. [FFGO23, Lemma E.1] gave a random $cd$-outcome POVM $\boldsymbol{P}_{d}$ on $\mathbb{C}^{d\times d}$ such that for all fixed density matrices $\rho\in\mathsf{D}(d)$, if $\boldsymbol{P}_{d}(\rho)$ denotes the probability distribution defined by performing $\boldsymbol{P}_{d}$ on $\rho$, then
@@
Now consider the case where ${\|\mathcal{M}-\mathcal{N}\|}_{J}\geq\varepsilon$. Let $\boldsymbol{\psi}\in\mathsf{D}(d_{\mathrm{in}})$ be the Haar random state from Algorithm 2, and let $\boldsymbol{X}=\mathopen{}\mathclose{{}\left\|\mathcal{M}(\boldsymbol{\psi})-I/d_{\mathrm{out}}}\right\|_{2}^{2}$. Fawzi et al. [FFGO23, middle of Page 40] proved that $\mathrm{Pr}\mathopen{}\mathclose{{}\left(\boldsymbol{X}\geq\operatorname*{\mathbb{E}}[\boldsymbol{X}]/2}\right)\geq 1/1000$. If $\boldsymbol{X}\geq\operatorname*{\mathbb{E}}[\boldsymbol{X}]/2$ and $v(\boldsymbol{P}_{d_{\mathrm{out}}}(\mathcal{M}(\boldsymbol{\psi})))\geq\sqrt{\boldsymbol{X}}/20$, an event which occurs with probability at least $1/2000$ by the definition of $\boldsymbol{P}_{d_{\mathrm{out}}}$, then
$\varepsilon$ $\leq{\|J_{\mathcal{M}}-I/d_{\mathrm{in}}d_{\mathrm{out}}\|}_{1}$
$\leq\sqrt{2(d_{\mathrm{in}}+1)d_{\mathrm{out}}\boldsymbol{X}}$
@@ 6.2 Upper bounds for erasure, unitary, and pure state replacement channels
Proof.
This is the case of Lemma B.2 where $\mathcal{L}=\mathcal{P}$ and $\mathcal{K}=\mathcal{I}$. ∎
Now we reduce the task of testing identity to $\mathcal{I}\otimes\mathcal{N}$ to that of testing identity to $\mathcal{N}$, for an arbitrary channel $\mathcal{N}$. To match the context in which we will apply this reduction, we phrase it in terms of ancilla-free, non-adaptive channel testers with perfect completeness, but the proof can easily be adopted to the other query models from Section 2.3 and to channel testers with imperfect completeness as well.
@@
and queries to $\mathcal{P}$ can be simulated using queries to $\mathcal{M}$. The identity channel is the erasure channel with input dimension equal to the output dimension, so the result follows by the above argument.
@@ 6.3 Lower bound for the completely depolarizing channel
We use the well-known fact that $d_{\mathrm{TV}}(P,Q)\leq\sqrt{2}\cdot d_{\mathrm{H}}(P,Q)$ for all distributions $P,Q$, where
$d_{\mathrm{H}}(P,Q)\coloneqq\sqrt{1-\sum_{x}\sqrt{P(x)Q(x)}}$
denotes Hellinger distance. Let $P=\bigotimes_{i=1}^{n}P_{i}$ and $Q=U^{\otimes n}$, and write
@@
which rearranges to $\sqrt{1-x}+\sqrt{1+x}\geq 2-x^{2}$. [^24] In fact, the stronger inequality $\sqrt{1-x}+\sqrt{1+x}\geq 2-(2-\sqrt{2})x^{2}$ holds, but proving this is more time-consuming and is not necessary for our purposes. It follows that
@@
Recalling 2.2 of ancilla-free, non-adaptive channel testers, write
@@ 6.5 Tomography
See 1.11
Proof.
O’Donnell and Wright [OW15, Theorem 1.10] gave an algorithm $\textsc{StateTom}_{\delta}$ that performs an entangled measurement on $O(d^{2}/\delta^{2})$ copies of a density matrix $\rho\in\mathbb{C}^{d\times d}$, and with probability at least $2/3$ outputs the description of a density matrix $\sigma\in\mathbb{C}^{d\times d}$ such that ${\|\rho-\sigma\|}_{1}\leq\delta$. Our algorithm is to first perform $\textsc{StateTom}_{\varepsilon/2}$ on $J_{\mathcal{M}}$, yielding the description of a state $\rho\in\mathsf{D}(d_{\mathrm{in}}\otimes d_{\mathrm{out}})$, and then output the description of a channel $\mathcal{N}$ that minimizes ${\|\rho-J_{\mathcal{N}}\|}_{1}$. If $\textsc{StateTom}_{\varepsilon/2}$ succeeds, then by the triangle inequality
${\|\mathcal{M}-\mathcal{N}\|}_{J}\leq{\|J_{\mathcal{M}}-\rho\|}_{1}+{\|\rho-J_{\mathcal{N}}\|}_{1}\leq 2{\|J_{\mathcal{M}}-\rho\|}_{1}\leq 2\cdot\varepsilon/2\leq\varepsilon.\qed$
@@
$\mathopen{}\mathclose{{}\left\|\sigma_{n}-\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[|\boldsymbol{\mathcal{N}}\rangle\!\langle\boldsymbol{\mathcal{N}}|^{\otimes 2}}\right]}\right\|_{1}=\mathopen{}\mathclose{{}\left\|\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[|\boldsymbol{\mathcal{N}}\rangle\!\langle\boldsymbol{\mathcal{N}}|\otimes\mathopen{}\mathclose{{}\left(\rho_{\boldsymbol{\mathcal{N}},n}-|\boldsymbol{\mathcal{N}}\rangle\!\langle\boldsymbol{\mathcal{N}}|}\right)}\right]}\right\|_{1}=\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left\|\rho_{\boldsymbol{\mathcal{N}},n}-|\boldsymbol{\mathcal{N}}\rangle\!\langle\boldsymbol{\mathcal{N}}|}\right\|_{1}\leq 0.2,$
@@
$\leq$ $\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[(S_{\boldsymbol{\mathcal{N}}}(\mathsf{B})+S_{\boldsymbol{\mathcal{N}}}(\mathsf{C}))-(S(\mathsf{C})-S(\mathsf{B}))-(S_{\boldsymbol{\mathcal{N}}}(\mathsf{C})-S_{\boldsymbol{\mathcal{N}}}(\mathsf{B}^{\prime}))+(S(\mathsf{B}^{\prime})+S(\mathsf{C}))}\right]$ Eqs. 11 and 12
@@ Acknowledgments
GR, AD, TG are supported by the EPSRC New Horizons grant EP/X018180/1. SS is supported by a Royal Commission for the Exhibition of 1851 Research Fellowship. TG and HA are supported by ERC Starting Grant 101163189 and UKRI Future Leaders Fellowship MR/X023583/1. We thank Min-Hsiu Hsieh, Tony Metger, Jon Wright, Henry Yuen, and Haimeng Zhao for helpful discussions.
@@ Appendix A Barriers to strengthening the results from Section 5
Appendix B Proof of Lemma 6.2
We use the following equality:
Lemma B.1.
For all matrices $X,Y\in\mathbb{C}^{d\times d}$,
where $\boldsymbol{\psi}\in\mathsf{D}(d)$ is Haar random.
Gu [Gu13] proved a significant generalization of Lemma B.1 using Weingarten calculus. For completeness and simplicity, below we present a self-contained proof of Lemma B.1 (essentially due to Montanaro and de Wolf [MdW13, proof of Proposition 21]) using only Eq. 4:
Proof.
We have
where the last equality is by Eq. 4. Clearly
and furthermore
The result follows by combining the above three equations. ∎
Lemma 6.2 is the case of the following where $\mathcal{L}=\mathcal{K}$ is the difference between two channels:
Lemma B.2.
where $\boldsymbol{\psi}\in\mathsf{D}(d)$ is Haar random.
Proof.